OpenStudy (anonymous):

Given the linear equations (i) 6x-3y=15 and (ii) 4x+10y-30=0 a) write the equation in y=mx+b form b)State the slope and y=intercept

OpenStudy (zzr0ck3r):

write each eqaution in that form?

OpenStudy (zzr0ck3r):

first one 6x-3y=15 -3y=15-6x y=2x-5 slope = 2 y int = -5 can you do number 2?

OpenStudy (anonymous):

Yes thank you :)

OpenStudy (radar):

It appears you must rearrange the equation getting the y variable on the left of the equal signt and the rest of it on the right of the equal sign.